### v1

@@ Abstract
We formulate and prove an Agreement Theorem for quantum mechanics (QM), describing when two agents, represented by separate laboratories, can or cannot maintain differing probability estimates of a shared quantum property of interest. Building on the classical framework (Aumann, 1976), we define the modality of “common certainty” through a hierarchy of certainty operators acting on each agent’s Hilbert space. In the commuting case – when all measurements and event projectors commute – common certainty leads to equality of the agents’ conditional probabilities, recovering a QM analog of the classical theorem. By contrast, when non-commuting operators are allowed, the certainty recursion can stabilize with different probabilities. This yields common certainty of disagreement (CCD) as a distinctive QM phenomenon. Agreement is restored once measurement outcomes are recorded in a classical register. The classical Agreement Theorem can therefore be seen as emergent from the quantum world via recording. We establish an impossibility result stating that QM forbids a scenario where one agent is certain that a property of interest occurs, and is also certain that the other agent is certain that the property does not occur. In this sense, QM admits non-classical disagreement, but disagreement is still bounded in a disciplined way. We argue that our analysis offers a rigorous approach to the longstanding issue of how to understand intersubjectivity across agents in QM.
@@
Soon after the formalism of quantum mechanics (QM) was in place, the question of the role of the observer arose. Bohr argued that the outcome of an experiment, and even which physical quantities can meaningfully be said to possess values, naturally depends on the experimental context, while Einstein famously objected that such “observer-created” facts conflict with scientific realism (Held, 1998). Wigner, (1961) sharpened the issue with his Friend Paradox. If Wigner treats his colleague and her laboratory as a quantum system in superposition, he must assign an entangled state to a situation in which the colleague herself has already witnessed a definite outcome. This thought experiment suggests that observer dependence might even yield intersubjective contradiction.
@@
The debate has continued. Elaborating on Wigner’s Friend, Frauchiger and Renner (2018) recently presented a scenario exhibiting an especially strong form of intersubjective disagreement, which causes the authors to question the consistency of QM. (See also Brukner, 2018; Nurgalieva and del Rio, 2018; Proietti et al., 2019). Under quantum Bayesianism (QBism), full intersubjectivity is allowed in any measurement setting (Fuchs and Schack, 2013; Fuchs, Mermin, and Schack, 2014). QBism is also interested in how agents might actively create conditions for agreement among them (Schack, 2024), such as when Alice first performs a measurement and can then predict with probability one the outcome that Bob will observe if he performs a measurement on her post-measurement state. In relational QM, every state of a system is only defined relative to another system – usually, an observer (Rovelli, 1996; Di Biagio and Rovelli, 2021). But in relational QM, too, the question of intersubjective agreement has been studied and a specific postulate, called “cross-perspective links” (Adlam and Rovelli, 2022), put forward that ensures agreement about past quantum events.
This paper aims to add to our understanding of intersubjectivity in QM by formally describing the epistemic states of agents, that is, what they know or believe about underlying properties of interest, know or believe about what other agents know or believe, and so on, potentially to arbitrarily many levels of reasoning. To do so, we draw on the well-established sub-field of game theory that formalizes the epistemic states of agents (Aumann, 1976; Aumann and Brandenburger, 1995; Dekel and Siniscalchi, 2015). Thus, our results operate at this intersection of the logic of belief and the physics of measurement in a quantum setting.
In our first scenario, Alice and Bob each perform local measurements in their respective labs and, updating via the Born-Lüders rule, obtain conditional probabilities for some property of shared interest. The probability estimates by Alice and Bob will generally differ, which is natural. Next suppose that Alice can calculate with certainty what Bob’s probability is. This number may be different from her own probability, which is again natural. Likewise with the roles of Alice and Bob interchanged. The issue of disagreement between Alice and Bob comes into sharp focus when we suppose that there is “common certainty” of their respective probabilities, that is, each agent is certain of the probability estimates, is certain the other agent is certain, and so on ad infinitum.
@@
We first study a commuting regime where Alice’s and Bob’s measurements commute with each other, and each commutes with the property of interest. Our first result is: In the commuting regime, if Alice and Bob have common certainty of their respective probability estimates of a shared property of interest $E$, then their estimates must agree. We then turn to a non-commuting regime, where an agent’s measurements do not necessarily commute with the property of interest. Our second result is: In a non-commuting regime, a quantum system can exhibit common certainty of disagreement (CCD). To show this, we produce a qutrit-qubit-qubit system in a particular entangled state where Alice’s measurements do not commute with the property of interest $E$, and it is common certainty that Alice assigns probability $1/2$ to $E$ and Bob assigns probability $1$ to $E$.
Our Commutative Agreement Theorem encompasses the fundamental Agreement Theorem of classical epistemics (Aumann, 1976), which states that if two agents share a common prior on a classical probability space, and their respective probability estimates of an event of interest are common knowledge, then these estimates must agree. (In the Supplementary Material, we explain the shift from knowledge to certainty modality, needed to treat quantum systems.) The Classical Agreement Theorem has been used, for example, to establish epistemic conditions for Nash equilibrium (Aumann and Brandenburger, 1995; Pacuit and Roy, 2025) and investigate market efficiency (Milgrom and Stokey, 1982; Sebenius and Geanakoplos, 1983; Gizatulina and Hellman, 2019).
We show that, in the non-commuting case, recording the outcomes of Alice’s and Bob’s measurements in a classical register restores agreement. This offers a new interpretation of the Classical Agreement Theorem as true even in a quantum world, under the physical act of classical recording. This is a different viewpoint from classical epistemics, which has not taken into account the physical environment in which agents operate.
Our final set of results provides bounds on disagreement. We establish a $0\operatorname{-}1$ impossibility result: There is no positive-probability state where Alice is certain of $E$ and certain that Bob is certain of not-$E$. By contrast, in the Supplementary Material, we produce a superquantum no-signaling box (Popescu and Rohrlich, 1994, that exhibits $0\operatorname{-}1$ disagreement. (These two results extend earlier work by Contreras-Tejada et al., 2021. By Abramsky and Brandenburger, 2011, the no-signaling example can also be cast in quasi-classical terms using signed probabilities, as we show.) Evidently, while QM permits “spooky” disagreement across agents in the form of CCD, it is not a case of anything goes: The extreme of $0\operatorname{-}1$ disagreement is ruled out. As for the possibility of $0\operatorname{-}1$ disagreement in no-signaling systems, we see this phenomenon as adding to the well-known list of implausible superquantum behaviors, such as trivialized communication complexity (van Dam, 2005), violations of information causality (Pawlowksi et al., 2009), and signaling in the classical limit (Rohrlich, 2014). Finally, we obtain two robustness results for the Commutative Agreement Theorem, showing smooth behavior under small departures from a shared initial state and exact common certainty (cf. Hellman, 2013; Monderer and Samet, 1989).
@@
We believe that our results help shed light on the issue of intersubjectivity in QM surveyed above. First, a clear distinction emerges between the commuting and non-commuting regimes: CCD is impossible in the former case and possible in the latter case. On the one hand, our findings challenge the notion that intersubjective disagreement should never be tolerated in a physical theory. The explicit construction of CCD shows that QM permits two agents to sustain different probability estimates of the same event, even under common certainty. We see this phenomenon as in line with the kind of perspective-dependent facts that QBism (Fuchs and Schack, 2013) and relational QM (Rovelli, 1996) were formulated to accommodate, and which more traditional interpretations of QM struggle to accept. Far from being a pathology, CCD appears as a natural consequence of non-commuting measurements and provides direct support for interpretations that treat quantum probabilities as agent-relative. In employing formal epistemics, our work may also suggest new tools for the further development of these interpretations.
On the other hand, we obtain bounds on the extent of disagreement that is possible. Our $0\operatorname{-}1$ impossibility theorem limits disagreement, even in the non-commuting case. So, we show that quantum theory is not entirely undisciplined and rules out certain epistemic scenarios.
@@
We next present our framework, the commuting scenario we analyze, and our first result on common certainty and agreement. We then present two examples – the first is an illustration of the commuting case, while the second is non-commuting and exhibits CCD. We then examine restoration of agreement via a classical register and bounds on disagreement. We conclude with a discussion of some interpretational aspects, related work, and possible future directions.
Agreement Theorem
@@
Alice and Bob perform projective measurements on their respective systems, in order to estimate the property of interest in the third system $\mathcal{H}_{C}$. That is, for each agent $X$ (Alice or Bob), there is a projective measurement $\Pi_{X}=\{P_{X}^{k}\}_{k=1}^{\ell_{X}}$ with projectors $P_{X}^{k}\in{\cal B(H})$. Each projector for Alice takes the form $P_{A}^{i}=\hat{P}_{A}^{i}\otimes\mathbb{I}_{B}\otimes\mathbb{I}_{C}$, for $i=1,\dots,\ell_{A}$, and each projector for Bob takes the form $P_{B}^{j}=\mathbb{I}_{A}\otimes\hat{P}_{B}^{j}\otimes\mathbb{I}_{C}$, for $j=1,\dots,\ell_{B}$, where each $\hat{P}^{k}_{X}\in{\cal B(H}_{X})$, and $\mathbb{I}_{X}$ and $\mathbb{I}_{C}$ are the identity operators. The property of interest $E$ in the inaccessible lab is modeled as a projector $P_{E}=\mathbb{I}_{A}\otimes\mathbb{I}_{B}\otimes\hat{P}_{E}\in{\cal B(H)}$ with $\hat{P}_{E}\in{\cal B(H}_{C})$. In this setting, each pair of projectors for Alice and Bob commute with each other:
@@
Next, fix a projector $P_{F}\in\cal{B(H)}$. The (conditional) probability that agent $X$ assigns to property $F$, given the outcome $P_{X}^{k}$ of the measurement, is given by the Born rule with a Lüders update:
@@
which, as a sum of mutually orthogonal projectors, is again a projector. The probability that agent $Y$ assigns to $Q_{X}(F)$, given the outcome $P_{Y}^{k}$, is $\operatorname{Prob}[Q_{X}(F);P_{Y}^{k}]$. If this probability is $1$, we say that agent $Y$ is certain that agent $X$ assigns probability $q$ to property $F$. A particular instance is $q=1$, in which case we say agent $Y$ is certain that agent $X$ is certain of property $F$. We also write $C_{X}(F)$ for $Q_{X}(F;1)$.
@@
(See the Supplementary Material for details.) By finiteness of the projective measurements $\Pi_{X}$, for $X=A,B$, we know there is a finite integer $N$ such that the recursion stabilizes: $A_{n}=A_{N}$ and $B_{n}=B_{N}$ for all $n\geq N$. Writing $A_{*}=A_{N}$ and $B_{*}=B_{N}$, we call $C_{*}=A_{*}\,B_{*}$ the common certainty projector. It is the fixed point of the certainty recursion, capturing the infinite iteration “Alice is certain that Bob is certain that Alice is certain $\dots$ that the property $E$ has probability $q_{A}$ (resp. $q_{B}$).” Since $A_{*}$ and $B_{*}$ commute, we can also write $C_{*}=B_{*}\,A_{*}$, a fact we will use later. If $q_{A}=q_{B}=1$, we simply say that there is common certainty of $E$. We now have the ingredients to define the epistemic condition of interest.
@@
The final step is to observe that because $P_{E}$ commutes with Alice’s measurement projectors, her probability estimate $q_{A}$ on her actual outcome $P_{A}^{i}$ agrees with the coarse-grained conditional on her common-certainty projector:
@@
Commuting and Non-Commuting Examples
To clarify the scope and implications of the our Commutative Agreement Theorem, we present two examples. The purpose of the first example, which satisfies the commutativity conditions of Equations 12, is to demonstrate that the certainty recursion does not necessarily stabilize immediately. This establishes that our first theorem has nontrivial epistemic content in that agents may need to refine what they are certain about. In the second example, Alice’s measurement does not commute with the property of interest, and we obtain common certainty of disagreement (CCD), in contrast with our first theorem.
Example 1: The global state is pure and entangled within each of Alice’s and Bob’s labs, with its entanglement extending to the third, inaccessible lab containing a single qubit. Alice and Bob each perform local rank-$2$ projective measurements on their respective subsystems. Specifically, Alice’s Hilbert space decomposes into two orthogonal two-dimensional subspaces and Bob’s Hilbert space into three orthogonal two-dimensional subspaces:
@@
Notice that the recursion does not stabilize immediately. (Variants of the example could be designed that require additional steps until stabilization.) Also, we obtain a range of examples between $q=1$ ($\theta=0$) and $q=0$ ($\theta=\pi$), showing that our Commutative Agreement Theorem is not restricted to edge cases.
@@
Therefore, choosing $q_{A}=1/2$ and $q_{B}=1$, we have $A_{0}=P_{A}^{0}+P_{A}^{1}$ and $B_{0}=P_{B}^{0}$. Moreover, $C_{A}(B_{0})=P_{A}^{0}+P_{A}^{1}$ and $C_{B}(A_{0})=P_{B}^{0}$, so that the recursion stabilizes immediately. Since $\operatorname{Tr}(A_{*}B_{*}\,\rho)=\operatorname{Tr}(B_{*}A_{*}\,\rho)=\operatorname{Tr}(P_{B}^{0}\,\rho)=2/3>0$, we conclude that it is common certainty that Alice holds the probability estimate $q_{A}=1/2$ and Bob holds the estimate $q_{B}=1$, in contrast with our Commutative Agreement Theorem. That is, we have exhibited a scenario of CCD. Later, we will establish a bound on the extent of disagreement possible in such scenarios.
It is instructive to see that Example 2 breaks only the final step (Equation 13) in the proof of the Commutative Agreement Theorem. It can be checked that Alice’s probability of $E$ conditioned on $A_{*}$ is $1$. But along her individual outcome branches, non-commutativity with $P_{E}$ reduces the conditional probability to $1/2$ (and to 0 on the third branch).
We chose a 33-22-22-level set-up in Example 2 because it naturally produces both CCD, as we have just seen, and also a branch exhibiting common certainty of agreement. For the latter, set $q_{A}=q_{B}=0$ and check that the corresponding certainty recursion stabilizes immediately with $\operatorname{Tr}(C_{*}\,\rho)=1/3>0$. It will be useful to have both branches when we revisit the example in the next section.
@@
In the full system, these operators lie in the commutative algebra ${\cal A}_{R}$ of Equation 23, and therefore:
@@
which says that the commutativity assumptions of Equations 1-2 are now satisfied on $\mathbb{I}_{\cal H}\otimes\mathcal{H}_{R}$.
For any projector $P\in\mathbb{I}_{\cal H}\otimes\mathcal{A}_{R}$ and any property $F$ acting on the system, the conditional probability of $F$ given $P$ is the analogous formula to Equation 3:
@@
where we use the recorded state $\rho^{\prime}$. With this definition, we can proceed to define analogs of the earlier certainty recursion of Equations 6-9. Letting $C_{*}^{\prime}$ be the stabilized common-certainty operator, our Commutative Agreement Theorem carries over directly.
@@
with nonzero blocks $(0,0)$, $(1,0)$, and $(2,1)$, each with weight $1/3$. The conditional probabilities of $E$, calculated from $\rho^{\prime}$, are:
@@
where $\operatorname{Tr}(C_{*}^{\prime}\,\rho^{\prime})=\frac{2}{3}>0$. Thus, there is recorded common certainty at $\rho^{\prime}$ that $q_{A}^{\prime}=q_{B}^{\prime}=\frac{1}{2}$. It can be checked that the other common-certainty branch in Example 2, with agreement $q_{A}=q_{B}=0$, is preserved under recording as a branch with $q_{A}^{\prime}=q_{B}^{\prime}=0$ and probability $\operatorname{Tr}(C_{*}^{\prime}\,\rho^{\prime})=\frac{1}{3}>0$.
@@
This result does not use any commutativity assumptions and therefore imposes an epistemic constraint independent of any compatibility conditions.
Our impossibility result has bite. The Supplementary Material contains an example of a no-signaling box (Popescu and Rohrlich, 1994) where 0-1 disagreement does arise: Alice is certain of $E$ and certain that Bob is certain of not-$E$. We also present an equivalent formulation of the example on phase space where the sole non-classicality is the use of signed probabilities (Abramsky and Brandenburger, 2011).

### v2

@@ Abstract
We formulate and prove an Agreement Theorem for quantum mechanics (QM), describing when two agents, represented by separate laboratories, can or cannot maintain differing probability estimates of a shared quantum property of interest. Building on the classical framework (Aumann, 1976), we define the modality of “common certainty” through a hierarchy of certainty operators acting on each agent’s Hilbert space. In the commuting case – when all measurements and event projectors commute – common certainty leads to equality of the agents’ conditional probabilities, recovering a QM analog of the classical theorem. By contrast, when non-commuting operators are allowed, the certainty recursion can stabilize with different probabilities. This yields common certainty of disagreement (CCD) as a distinctive QM phenomenon. We show that agreement will nevertheless re-emerge if measurement outcomes are recorded in a classical register. We also establish an impossibility result stating that QM forbids a scenario where one agent is certain that a property of interest occurs, and is also certain that the other agent is certain that the property does not occur. In this sense, QM admits non-classical disagreement, but disagreement is still bounded in a disciplined way. We argue that our analysis offers a rigorous approach to the longstanding issue of how to understand intersubjectivity across agents in QM.
1 Introduction
@@
Soon after the formalism of quantum mechanics (QM) was in place, the question of the role of the observer arose. Bohr argued that the outcome of an experiment, and even which physical quantities can meaningfully be said to possess values, naturally depend on the experimental context, while Einstein famously objected that such “observer-created” facts conflict with scientific realism (Held, 1998). Wigner (1961) sharpened the issue with his Friend paradox. If Wigner treats his colleague together with her laboratory as a quantum system in superposition, he must assign an entangled state to a situation in which the colleague herself has already witnessed a definite outcome. This thought experiment suggests that observer dependence might even yield intersubjective contradiction.
@@
The debate has continued in modern form. Elaborating on Wigner’s Friend, Frauchiger and Renner (2018) presented a scenario exhibiting an especially strong form of intersubjective disagreement, raising questions about the consistency of QM. (See also Brukner, 2018; Nurgalieva and del Rio, 2018; Proietti et al., 2019). Relational quantum mechanics (Rovelli, 1996; Di Biagio and Rovelli, 2021) holds that physical states are always defined relative to another system – usually, an observer. The question of intersubjective agreement has been studied under this approach and a specific postulate, called “cross-perspective links” (Adlam and Rovelli, 2023), put forward that ensures agreement about past quantum events. In quantum Bayesianism (QBism), full intersubjectivity is allowed in any measurement setting (Fuchs and Schack, 2013; Fuchs, Mermin, and Schack, 2014). QBism, too, is interested in how agents might actively create conditions for agreement among them (Schack, 2024), such as when Alice first performs a measurement and can then predict with probability one the outcome that Bob will observe if he performs a measurement on her post-measurement state.
Across these approaches, a basic question persists: When, if ever, can two agents sustain differing probability assignments about the same physical property of a quantum system, when they engage in reasoning about each other’s reasoning about these probability assignments?
Independently, the epistemic approach to game theory has developed a rigorous formal language for describing what agents believe, what they believe about one another’s beliefs, and so on to arbitrarily many levels (Aumann, 1976; Aumann and Brandenburger, 1995; Dekel and Siniscalchi, 2015). Within this framework, the classical Agreement Theorem (Aumann, 1976) shows that if two agents share a common prior and it is common knowledge (or common certainty) what probabilities they assign to an event, then those probabilities must agree. In slogan form, rational agents cannot “agree to disagree.” This result has been used, for example, to analyze Nash equilibrium (Aumann and Brandenburger, 1995; Pacuit and Roy, 2025) and market efficiency (Milgrom and Stokey, 1982; Sebenius and Geanakoplos, 1983; Gizatulina and Hellman, 2019).
In this paper, we bring these traditions together. We formulate and prove several theorems on agreement and disagreement when Alice and Bob are observers of a quantum system. Our analysis thereby operates at the intersection of the logic of belief and the physics of measurement. Concretely, we consider a setting in which Alice and Bob perform local measurements in their respective laboratories on subsystems of a shared quantum state and are both interested in estimating a property of a third subsystem. Each agent updates their own probability for the property of interest using the Born-Lüders rule. We then formalize a quantum version of “common certainty,” under which each agent is certain of the other agent’s probability estimate, is certain the other agent is certain of the first agent’s estimate, and so on ad infinitum. This is the epistemic condition we use to formalize intersubjectivity in the quantum context. Our first two results establish a clear structural distinction between commuting and non-commuting regimes.
@@
Commutative Agreement Theorem: In the commuting regime where Alice’s and Bob’s measurements commute with one another and with the property of interest, if Alice and Bob have common certainty of their respective probability estimates of that property, then their estimates must agree.
This result provides a direct quantum analog of the classical Agreement Theorem and shows that classical intersubjective agreement emerges naturally in QM whenever the relevant operators are mutually compatible.
Non-Commutative Disagreement Theorem: In a non-commuting regime where an agent’s measurement does not commute with the property of interest, it is possible that Alice and Bob have common certainty of their respective probability estimates, yet those estimates differ.
To establish this result, we produce a qutrit-qubit-qubit system in a particular entanglement state where Alice’s measurement does not commute with the property of interest, and it is common certainty that Alice assigns a probability $1/2$ to the property of interest while Bob assigns a probability of $1$ to this same property. We call this a situation of common certainty of disagreement (CCD). It is evidently a distinctive quantum phenomenon that cannot arise classically. Our third result establishes a limit on such epistemic conflict in QM.
$0\operatorname{-}1$ Impossibility Theorem: There is no positive-probability state in a (finite-dimensional) quantum system where Alice is certain of a property and is also certain that Bob is certain of the negation of that property.
This result validates that our epistemic analysis has physical bite – there are epistemic states ruled out by QM. It also distinguishes our approach from Frauchiger and Renner (2018), who, in a challenge to the internal consistency of QM, produce the “paradoxical” $0\operatorname{-}1$ scenario by chaining together statements of Alice’s certainty across mutually incompatible measurement branches. We show that this scenario cannot arise at a single state. However, we do produce a superquantum no-signaling box (Popescu and Rohrlich, 1994) that exhibits $0\operatorname{-}1$ disagreement. (By Abramsky and Brandenburger, 2011, this example can also be cast in quasi-classical terms using signed probabilities, as we show.) We see this phenomenon of extreme disagreement as adding to the well-known list of implausible superquantum behaviors, such as trivialized communication complexity (van Dam, 2005), violations of information causality (Pawlowksi et al., 2009), and signaling in the classical limit (Rohrlich, 2014).
Beyond these main structural results, we show how classical agreement can be recovered even in a non-commuting regime through the physical act of recording measurement outcomes in a classical register. This offers a new perspective on the classical Agreement Theorem, which can be understood as emergent from quantum theory via classical recording. We also establish two robustness results for our Commutative Agreement Theorem. First, if Alice and Bob start with different initial states, as might happen because of noise, channel disturbance, or the presence of an eavesdropper, the difference between their probability estimates can be quantitatively bounded (cf. Hellman, 2013). Second, if certainty is relaxed to approximate certainty, reflecting finite statistics, detector inefficiencies, or coarse-graining, then disagreement is again bounded (cf. Monderer and Samet, 1989). These results show that agreement in the commuting regime is stable under realistic perturbations.
@@
Taken together, our findings clarify the structure of intersubjectivity in QM. In the commuting regime, classical-style agreement prevails. In the non-commuting regime, CCD becomes possible. We therefore identify a clear boundary between classical and commuting intersubjectivity (ruled out under common certainty) and non-commuting intersubjectivity (allowed under common certainty). This second result, in particular, challenges the notion that intersubjective disagreement should never be tolerated in a physical theory. We see this finding as broadly in line with the kind of perspective-dependent facts that relational QM (Rovelli, 1996) and QBism (Fuchs and Schack, 2013) were formulated to accommodate. Far from being a pathology, CCD appears as a natural consequence of non-commuting measurements and provides direct support for interpretations that treat QM probabilities as agent-relative. In employing formal epistemics, our work may suggest new tools for further development of these interpretations.
@@
We next present our framework and establish our three main results. We then examine restoration of agreement via a classical register and bounds on disagreement. We conclude with a discussion of some interpretational aspects, related work, and possible future directions.
Commutative Agreement Theorem
@@
Alice and Bob perform projective measurements on their respective systems, in order to estimate the property of interest in the third system $\mathcal{H}_{C}$. That is, for each agent $X$ (Alice or Bob), there is a projective measurement $\Pi_{X}=\{P_{X}^{k}\}_{k=1}^{\ell_{X}}$ with projectors $P_{X}^{k}\in{\cal B(H})$. Each projector for Alice takes the form $P_{A}^{i}=\hat{P}_{A}^{i}\otimes\mathbb{I}_{B}\otimes\mathbb{I}_{C}$, for $i=1,\dots,\ell_{A}$, and each projector for Bob takes the form $P_{B}^{j}=\mathbb{I}_{A}\otimes\hat{P}_{B}^{j}\otimes\mathbb{I}_{C}$, for $j=1,\dots,\ell_{B}$, where each $\hat{P}^{k}_{X}\in{\cal B(H}_{X})$, and $\mathbb{I}_{X}$ and $\mathbb{I}_{C}$ are the identity operators. (We comment later on extension to a POVM framework.) The property of interest $E$ in the inaccessible lab is modeled as a projector $P_{E}=\mathbb{I}_{A}\otimes\mathbb{I}_{B}\otimes\hat{P}_{E}\in{\cal B(H)}$ with $\hat{P}_{E}\in{\cal B(H}_{C})$. In this setting, each pair of projectors for Alice and Bob commute with each other:
@@
Next, fix a projector $P_{F}\in\cal{B(H)}$. The (conditional) probability that agent $X$ assigns to property $F$, given the outcome $P_{X}^{k}$ of their measurement, is given by the Born rule with a Lüders update:
@@
which, as a sum of mutually orthogonal projectors, is again a projector. The probability that agent $Y$ assigns to $Q_{X}(F;q)$, given their outcome $P_{Y}^{k}$, is $\operatorname{Prob}[Q_{X}(F;q);P_{Y}^{k}]$. If this probability is $1$, we say that agent $Y$ is certain that agent $X$ assigns probability $q$ to property $F$. A particular instance is $q=1$, in which case we say agent $Y$ is certain that agent $X$ is certain of property $F$. We also write $C_{X}(F)$ for $Q_{X}(F;1)$.
@@
(See the Supplementary Material for details.) By finiteness of the projective measurements $\Pi_{X}$, for $X=A,B$, there is a finite integer $N$ such that the recursion stabilizes: $A_{n}=A_{N}$ and $B_{n}=B_{N}$ for all $n\geq N$. Writing $A_{*}=A_{N}$ and $B_{*}=B_{N}$, we call $C_{*}=A_{*}\,B_{*}$ the common certainty projector. It is the fixed point of the certainty recursion, capturing the infinite iteration “Alice is certain that Bob is certain that Alice is certain $\dots$ that the property $E$ has probability $q_{A}$ (resp. $q_{B}$).” Since $A_{*}$ and $B_{*}$ commute, we can also write $C_{*}=B_{*}\,A_{*}$, a fact we will use later. If $q_{A}=q_{B}=1$, we simply say that there is common certainty of $E$. We now have the ingredients to define the epistemic condition of interest.
@@
The final step is to observe that because $P_{E}$ commutes with Alice’s measurement projectors, her probability estimate $q_{A}$ when she measures and then obtains the outcome $P_{A}^{i}$ agrees with the coarse-grained conditional on her common-certainty projector:
@@
We now provide an example to show that the certainty recursion does not necessarily stabilize immediately. This establishes that our commutative theorem has nontrivial epistemic content in that agents may need to refine what they are certain about.
Example 1: Alice and Bob each perform local rank-$2$ projective measurements on their respective subsystems. Specifically, Alice’s Hilbert space decomposes into two orthogonal two-dimensional subspaces and Bob’s Hilbert space into three orthogonal two-dimensional subspaces:
@@
We see in this example that the recursion does not stabilize immediately. (Variants of the example could be designed that require additional steps until stabilization.) Also, we obtain a range of examples between $q=1$ ($\theta=0$) and $q=0$ ($\theta=\pi$), showing that our Commutative Agreement Theorem is not restricted to edge cases.
Non-Commutative Disagreement Theorem
The setup is exactly the same as in the previous section, except that we no longer require all the commutativity conditions of Equations 1 and 2 to hold. This breaks the Agreement Theorem and allows the distinctive phenomenon of common certainty of disagreement (CCD) to arise.
Non-Commutative Disagreement Theorem: In the non-commuting regime, it can be common certainty at state $\rho$ that Alice assigns $q_{A}$ to $E$ and Bob assigns $q_{B}$ to $E$, while $q_{A}\not=q_{B}$.
This result is established via the next example. (Later, we will establish a bound on the extent of disagreement possible in such scenarios.)
@@
Therefore, choosing $q_{A}=1/2$ and $q_{B}=1$, we have $A_{0}=P_{A}^{0}+P_{A}^{1}$ and $B_{0}=P_{B}^{0}$. Moreover, $C_{A}(B_{0})=P_{A}^{0}+P_{A}^{1}$ and $C_{B}(A_{0})=P_{B}^{0}$, so that the recursion stabilizes immediately. Since $\operatorname{Tr}(A_{*}B_{*}\,\rho)=\operatorname{Tr}(B_{*}A_{*}\,\rho)=\operatorname{Tr}(P_{B}^{0}\,\rho)=2/3>0$, we conclude that it is common certainty that Alice holds the probability estimate $q_{A}=1/2$ and Bob holds the estimate $q_{B}=1$. We have actually exhibited a family of scenarios of CCD, since the phenomenon does not depend on Bob’s measurement basis.
Note that in our non-commutative theorem, we do not include in the agents’ information states any variable recording measurement order or stage of protocol execution, nor any higher-order beliefs about such facts. Bob’s probability assignment is obtained by conditioning the common prior $\rho$ only on his own outcome. By no-signaling, and absent any public protocol information or other classical side information, Bob cannot infer from his local quantum data whether Alice has already measured. One could imagine a richer model in which Bob assigns probabilities to such protocol facts, to Alice’s beliefs about them, to Bob’s beliefs about Alice’s beliefs, and so on. The most natural interpretation of our Non-Commutative Disagreement Theorem is that it is “protocol-agnostic.” In an elaborated model with explicit order uncertainty, one may view it as the special case in which it is common certainty that Bob assesses $E$ before Alice’s measurement is incorporated. In the Discussion section, we consider the possibility of extending the protocol-agnostic framework to an epistemic model in which measurement order is an additional object of uncertainty.
@@
We now move from the protocol-agnostic setting to a protocol-explicit one. Specifically, we suppose that the agents’ measurement outcomes are recorded in a classical register encoding the transcripts of a shared recording protocol. This is not merely a notational refinement. Once such a register is included, facts about which measurements have been performed and which outcomes have occurred become part of the physical state itself, and the agents may condition on them through a commuting algebra. We show that this added classical structure removes the mechanism behind the common-certainty-of-disagreement phenomenon exhibited above and restores agreement. In this sense, the classical Agreement Theorem reappears as an emergent consequence of quantum theory once measurement histories are classically recorded.
@@
In the full system, these operators lie in the commutative algebra ${\cal A}_{R}$ of Equation 23, and therefore, for all pairs of classical outcomes $(i,j)$:
@@
which says that the commutativity assumptions of Equations 12 are now satisfied on $\mathbb{I}_{\cal H}\otimes\mathcal{H}_{R}$.
For any projector $P\in\mathcal{A}_{R}$ and any property $F$ acting on the system, the conditional probability of $F$ given $P$ is the analogous formula to Equation 3:
@@
where we use the recorded state $\rho^{\prime}$. With this definition, we can proceed to define analogs of the earlier certainty recursion of Equations 69. Letting $C_{*}^{\prime}$ be the stabilized common-certainty operator, our Commutative Agreement Theorem carries over directly.
@@
with nonzero blocks $(0,0)$, $(1,0)$, and $(2,1)$, each with trace $1/3$. The conditional probabilities of $E$, calculated from $\rho^{\prime}$, are:
@@
where $\operatorname{Tr}(C_{*}^{\prime}\,\rho^{\prime})=\frac{2}{3}>0$.
Thus, there is recorded common certainty at $\rho^{\prime}$ that $q_{A}^{\prime}=q_{B}^{\prime}=\frac{1}{2}$. It can be checked that the other common-certainty branch in Example 2, with agreement $q_{A}=q_{B}=0$, is preserved under recording as a branch with $q_{A}^{\prime}=q_{B}^{\prime}=0$ and probability $\operatorname{Tr}(C_{*}^{\prime}\,\rho^{\prime})=\frac{1}{3}>0$.
@@
This result does not use any commutativity assumptions and therefore imposes an epistemic constraint independent of any compatibility conditions. Our impossibility theorem has bite. The Supplementary Material contains an example of a no-signaling box (Popescu and Rohrlich, 1994) where 0-1 disagreement does arise: Alice is certain of $E$ and certain that Bob is certain of not-$E$. We also present an equivalent formulation of the example on phase space where the sole non-classicality is the use of signed probabilities (Abramsky and Brandenburger, 2011).